OpenStudy (anisah.21a):

2x=50-y solve for either or

OpenStudy (madhu.mukherjee.946):

x=25-y/2

OpenStudy (anisah.21a):

can you show your work please and thanks!

OpenStudy (madhu.mukherjee.946):

2x=50-y now taking two on the other side and dividing the expression by two x=25-y/2